OpenStudy (anonymous):

Let f(x)=-2x-7 and G(x)=-4x+6. Find (g o f)(-5)

OpenStudy (austinl):

Ok, we have: \(f(x)=-2x-7\) \(g(x)=-4x+6\) We want \((gof)(x)=g(f(x))\) So that means that we plug f(x) into g(x) wherever there is an x. \(g(f(x))=-4(-2x-7)+6\rightarrow 8x+34\)

OpenStudy (anonymous):

The answers that I can have are A)26 B)3 C)-59 D)-6

OpenStudy (austinl):

I was making sure that you were following me.

OpenStudy (anonymous):

I follow you.

OpenStudy (anonymous):

I got to that point but I can't figure out how to get a whole number answer.

OpenStudy (austinl):

Then you plug in - for all instances of x :)

OpenStudy (anonymous):

Plug in negative?

OpenStudy (austinl):

-5* My numlock was off...

OpenStudy (anonymous):

Thanks a ton! I got it! It was -6!

OpenStudy (austinl):

No problemo!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."